OpenStudy (anonymous):

Josiah traveled 196 miles on 8 gallons of gasoline. At this rate, how many gallons of gasoline would he need to travel 343 miles? a. 9 c. 21 b. 14 d. 16.5

OpenStudy (anonymous):

@Eoir

OpenStudy (anonymous):

please help i am so tired i cant even comprehend what oi am reading please just help me get threw this 10 questiin qui

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

@Desireelover

OpenStudy (anonymous):

@rockstar0765

OpenStudy (anonymous):

@hughfuve

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@nikvist

OpenStudy (anonymous):

please

OpenStudy (kropot72):

He can travel 196/8 miles on 1 gallon. Therefore the number of gallons needed to travel 343 miles is given by: \[\large \frac{343}{\frac{196}{8}}=\frac{343\times8}{196}=you\ can\ calculate\]

OpenStudy (anonymous):

thanks

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

14 correct?

OpenStudy (anonymous):

i am very tired so please bare with me if I ma rong